OpenStudy (anonymous):

what the lcm of 5 and 9

OpenStudy (anonymous):

45

myininaya (myininaya):

5(9)

jimthompson5910 (jim_thompson5910):

5 and 9 have no common factors, so the LCM is just the product of 5 and 9 So the LCM is 5*9 = 45

OpenStudy (anonymous):

LCm= 5*9 =45

OpenStudy (saifoo.khan):